# The Baryonic Tully-Fisher Relation of HI-bearing Low Surface Brightness Galaxies Implies Their Formation Mechanism

### v2

@@ 6 Summary
Taking advantage of the ALFALFA HI survey data, we have selected the HI-bearing LSBGs and HSBGs, and estimated their baryonic masses $M_{\rm b}$ and rotation velocities $v_{\rm rot}$. Our findings reveal that, statistically, the HI-bearing LSBGs and HSBGs also conform to the BTFR found in typical late-type galaxies as well as ordinary dwarf galaxies with slope of $\beta\sim 4$. These results hint that LSBGs may owe their properties to the high spins of their host dark matter halos, rather than the halo densities or feedback processes. Additionally, our findings do not rule out the Modified Newtonian Dynamics (MOND) theory.
Our findings conflict with the BTFR results of UDGs (Karunakaran et al., 2020; Hu et al., 2023; Rong et al., 2024). This is primarily because that our LSBG sample contains very few UDGs. Indeed, almost all of our LSBGs are classic LSBGs, rather than UDGs with extremely large effective radii. Therefore, we propose that UDGs may be a distinct galaxy population with different BTFR and mass distributions, compared with typical LSBGs. They may also have distinct formation mechanisms that warrant further investigations.
@@ Acknowledgements
We sincerely thank Xufen Wu, Haochen Jiang, Jingshuo Yang, Shijiang Chen and Yu Chen for the inspiring discussions during this study. Y.R. acknowledges supports from the NSFC grant 12273037, and the CAS Pioneer Hundred Talents Program (Category B), as well as the USTC Research Funds of the Double First-Class Initiative (No. YD2030002013). H.H. is supported by the Fundamental Research Funds for the Central Universities, the CAS Project for Young Scientists in Basic Research Grant No. YSBR-062, the National SKA Program of China No. 2022SKA0110201, and the NSFC grant No. 12033008.

### v3

@@ 6 Summary
Taking advantage of the ALFALFA HI survey data, we have selected the HI-bearing LSBGs and HSBGs, and estimated their baryonic masses $M_{\rm b}$ and rotation velocities $v_{\rm rot}$. Our findings reveal that, statistically, the HI-bearing LSBGs and HSBGs also conform to the BTFR found in typical late-type galaxies as well as ordinary dwarf galaxies with slope $\sim 4$. These results hint that LSBGs may owe their properties to the high spins of their host dark matter halos, rather than the halo densities or feedback processes. Additionally, our findings do not rule out the Modified Newtonian Dynamics (MOND) theory.
Our findings conflict with the BTFR results of UDGs (Mancera Piña et al., 2019; Guo et al., 2020; Hu et al., 2023; Karunakaran et al., 2020; Rong et al., 2024). This is primarily because that our LSBG sample contains very few UDGs. Indeed, almost all of our LSBGs are classic LSBGs, rather than UDGs with extremely large effective radii. Therefore, we propose that UDGs may be a distinct galaxy population with different BTFR and mass distributions, compared with classical LSBGs. They may also have distinct formation mechanisms that warrant further investigations.
@@ Acknowledgements
We express our sincere thanks to the referee for the detailed comments and suggestions. We also thank Xufen Wu, Haochen Jiang, Jingshuo Yang, Shijiang Chen and Yu Chen for the inspiring discussions during this study. Y.R. acknowledges supports from the NSFC grant 12273037, the CAS Pioneer Hundred Talents Program (Category B), the USTC Research Funds of the Double First-Class Initiative (No. YD2030002013), and the research grants from the China Manned Space Project (the second-stage CSST science projects: “Investigation of small-scale structures in galaxies and forecasting of observations” and “CSST study on specialized galaxies in ultraviolet and multi-band”). H.H. is supported by the Fundamental Research Funds for the Central Universities, the CAS Project for Young Scientists in Basic Research Grant No. YSBR-062, the National SKA Program of China No. 2022SKA0110201, and the NSFC grant No. 12033008.